OpenStudy (anonymous):

find range and domain of (x^2 + 2x + 1)/(x^2-8x+12)?

OpenStudy (anonymous):

Can you factorize denominator ??

OpenStudy (anonymous):

\[x^2 - 8x + 12\neq0\]

OpenStudy (anonymous):

Yes.. Firstly factorize it..

OpenStudy (anonymous):

(x-6) (x-2)

OpenStudy (anonymous):

Yes So what x cannot take here?

OpenStudy (anonymous):

so Domain = R-{ 2 ,6}

OpenStudy (anonymous):

Yes right..

OpenStudy (anonymous):

Then range

OpenStudy (anonymous):

For range put f(x) = y firstly factorize the numerator here..

OpenStudy (anonymous):

so.......wat to do?

OpenStudy (anonymous):

Call @experimentX here.

OpenStudy (anonymous):

I am bad experience when quadratic comes in range..

OpenStudy (experimentx):

can you make partial fractions?

OpenStudy (anonymous):

wat is partial fractions?

OpenStudy (experimentx):

wait ... i'll look for a method.

OpenStudy (anonymous):

y = f(x) y(x^2-8x+12) = x^2 + 2x +1

OpenStudy (anonymous):

x^2(y-1) - x(8y + 2) + (12y -1) =0

OpenStudy (anonymous):

i got such quad eq

OpenStudy (anonymous):

i got 4y^2 + 21 >= 0

OpenStudy (experimentx):